# Advancing Multi-Modal Sensing through Expandable Modality Alignment

### v1

@@ Abstract.
Sensing technology is widely used for comprehending the physical world, with numerous modalities explored in past decades. While there has been considerable work on multi-modality learning, they all require data of all modalities be paired. How to leverage multi-modality data with partially pairings remains an open problem.
To tackle this challenge, we introduce the Babel framework, encompassing the neural network architecture, data preparation and processing, as well as the training strategies. Babel serves as a scalable pre-trained multi-modal sensing neural network, currently aligning six sensing modalities, namely Wi-Fi, mmWave, IMU, LiDAR, video, and depth. To overcome the scarcity of complete paired data, the key idea of Babel involves transforming the N-modality alignment into a series of two-modality alignments by devising the expandable network architecture. This concept is also realized via a series of novel techniques, including the pre-trained modality tower that capitalizes on available single-modal networks, and the adaptive training strategy balancing the contribution of the newly incorporated modality with the previously established modality alignment.
Evaluation demonstrates Babel’s outstanding performance on eight human activity recognition datasets, compared to various baselines e.g., the top multi-modal sensing framework, single-modal sensing networks, and multi-modal large language models. Babel not only effectively fuses multiple available modalities (up to 22% accuracy increase), but also enhance the performance of individual modality (12% averaged accuracy improvement). Case studies also highlight exciting application scenarios empowered by Babel, including cross-modality retrieval (i.e., sensing imaging), and bridging LLM for sensing comprehension.
The Work is done during internship at Microsoft Research
copyright: acmcopyrightjournalyear: 2023copyright: rightsretainedconference: Preprint; price: ;
@@ 1. Introduction
Motivation: Sensing offers unique abilities to perceive the physical world. It has been widely deployed in various applications across diverse fields, including health-care, mixed reality, smart driving, and many others. Over the past decades, numerous sensing modalities have been explored (Pu et al., 2013; Adib and Katabi, 2013; Zhao et al., 2018b; Wang et al., 2014; Wei and Zhang, 2016, 2015; mmM, 2021; Nandakumar et al., 2015; Mao et al., 2016). Each sensing modality provides a unique and complementary viewpoint for observing the world, thereby necessitating the simultaneous use of multiple sensing modalities, known as multi-modal sensing.
@@
Early methods for organizing multiple sensing modalities relied on handcrafted heuristics or features (Zhou et al., 2014), which is proved difficult to scale across various tasks due to the complexity of sensing signals and environments. Recent advancements in Deep Learning (DL) based methods have offered promising solutions (Zhao et al., 2018a; Deldari et al., 2022; Jain et al., 2022). Given paired modality data as inputs, DL methods could identify the complementarity among various sensing modalities, a process known as modality alignment.
Figure 1. Align multiple sensing modalities into one unified representation to enhance sensing and empower new applications.
Modality alignment projects the representations of each sensing modality into a unified and shared space, as depicted in Fig. 1. The alignment process could empower sensing in several ways: firstly, the alignment process restructures the features of one modality with the help from another, thereby allowing modalities to complement each other. Secondly, the unified feature space enables the ease of fusing multiple modalities. More intriguingly, the aligned modalities could potentially give rise to new applications. One could utilize the unified feature space as the proxy, using the signals from one sensing modality to retrieve the representations of another modality (e.g., leveraging Wi-Fi channel state information (CSI) to obtain visual representation, thus enabling Wi-Fi imaging). The aligned features across modalities could also naturally serve as a protocol to bridge sensing abilities with Large Language Models (LLMs).
To harness the power of modality alignment, our goal is to develop a pre-trained foundation sensing model that is capable of aligning a wide range of prevalent sensing modalities. This would enable the employment of either a singular modality or a fusion of aligned modalities for downstream sensing tasks.
@@
Challenges: Although modality alignment in AI is a growing research area, its application in sensing presents significant hurdles. The fundamental challenge in supporting multi-modality in wireless sensing is the data scarcity, specifically, (i) the scarcity of paired data, which is essential for aligning two modalities, and (ii) the scarcity of multi-paired modalities. For instance, the widely-used CLIP (Radford et al., 2021) required 400 million image-text pairs for pre-training. In sensing, there lacks paired data from all modalities since some modality data require specialized hardware and expertise to collect. Existing datasets only contain data from a subset of modalities (e.g., up to 3) (Chen et al., 2015; Yang et al., 2023c; Bocus et al., 2021; Kay et al., 2017). How to take advantage of partially paired data becomes increasingly important.
Despite considerable work, existing research (Han et al., 2023; Yuan et al., 2024; Ouyang et al., 2022; Xu et al., 2023; Zhang et al., 2023a; Girdhar et al., 2023) struggle to fully incorporate multiple sensing modalities. For instance, due to the shortage of paired data, OneLLM (Han et al., 2023) supports a limited number of sensing modalities, i.e., IMU only, with subpar performance (see Table 4). Cosmos (Ouyang et al., 2022) pioneered the alignment of multiple modalities, but due to the scarcity of multi-paired modalities, it aligns a limited number of modalities e.g., IMU and depth.
@@
Our Work: We present Babel, establishing the first scalable pre-trained network aligning six sensing modalities as shown in Fig. 1. The design of Babel is underpinned by two key insights. (i) Despite the scarcity of paired data, there exist well-developed encoders or feature extractors for single modality sensing, which have been extensively explored by experts. By leveraging these encoders, the amount of paired data required for modality alignment can be significantly reduced. (ii) Even though few datasets provide more than three paired modalities, numerous paired datasets exist that share common sensing modalities. These shared modalities can serve as a bridge for multi-modality alignment (see Fig. 2).
Drawing from these insights, the key idea of Babel is to achieve expandable multi-sensing modality alignment by transforming an N-modality alignment problem into a sequence of binary modality alignments. The expandability presents a dual advantage: firstly, it allows for the integration of a new modality by aligning it with a previously incorporated modality in the network; secondly, it offers the capability to adapt to each different dataset by absorbing new insights while preserving the knowledge of previously aligned modalities. This expandability facilitates the effective utilization of partially paired data in the sensing community.
@@
To realize this expandability, we innovatively introduce three key techniques: the pre-trained modality tower (§4), the expandable network architecture (§5), and the adaptive training strategy (§6). Each modality utilizes a modality tower to extract features from raw data. We build these towers using existing singular-modal sensing feature extractors (e.g., LIMU-BERT (Xu et al., 2022) for IMU), and extent with concept alignment modules for the contrastive learning with other modality towers. The expandable network architecture enables sequential training phases with only paired samples. Within it we also propose the prototype network, shared by all modalities, maintains the knowledge of aligned modalities when adding new ones. Lastly, our adaptive training strategy balances the contribution of newly added modalities to the unified representation, optimally assimilating new knowledge during model growth without disrupting established alignments.
@@
We offer a comprehensive implementation of Babel, encompassing the network architecture, data preparation and processing, as well as the training strategies. In Babel, we currently align six sensing modalities: two for wireless sensing, namely Wi-Fi and mmWave, two for mobile sensing, specifically IMU and LiDAR, and two for general vision, namely RGB and depth. As an expandable framework, Babel is allowed for aligning more modalities in the future by the community effort. Aligned modalities can be chosen for downstream tasks, either individually or in combination. Eventually five datasets are utilized to construct Babel, including UTD-MHAD (Chen et al., 2015), Kinetics-400 (Kay et al., 2017), OPERANet (Bocus et al., 2021), XRF55 (Wang et al., 2024), and MM-Fi (Yang et al., 2023c).
The current pre-trained Babel is evaluated on the typical downstream sensing task of Human Activity Recognition (HAR), across eight datasets, which include both multi-modal and singular-modal datasets (Chen et al., 2015; Yang et al., 2023c; Bocus et al., 2021; Wang et al., 2024; Reyes-Ortiz et al., 2016; Zhang et al., 2022; An et al., 2022a; Li et al., 2010). To demonstrate Babel’s capability in modality enhancement and fusion, we compared it to an array of baselines, including the state-of-the-art (SOTA) multi-modal sensing framework (Ouyang et al., 2022), singular-modal sensing networks (Xu et al., 2021; Yang et al., 2023a; Li et al., 2010; An et al., 2022b), and the emerging Multi-modal Large Language Models (MLLMs) (Han et al., 2023; Yuan et al., 2024; Girdhar et al., 2023; Zhang et al., 2023a). Specifically, owing to the alignment across numerous sensing modalities, Babel improves the accuracy by up to 20%, and 12% on average across six modalities, compared to the performance before the alignment. Compared to the SOTA singular-modal network, Babel also brings the consistent accuracy improvements across various datasets. Due to the unified representation space, Babel increases multi-modal sensing fusion accuracy by up to 22% compared to current multi-modal frameworks. When comparing with MLLMs, Babel surpasses them by 25.2% across HAR datasets. In additional to HAR, we also present two real application case studies to highlight Babel’s potential. The first is sensing imaging to illustrate cross-modality retrieval. With Babel, the original image-to-image unCLIP (Ramesh et al., 2022) diffusion model can be supplemented with non-visual data as input to generate images. The other case aims to bridge the gap between LLM and sensing. By injecting the IMU sensing signal through Babel into the Video-LLaMA (Zhang et al., 2023b), the LLM can understand the sensing signals without any training.
@@
  • Babel, to the best of our knowledge, is the first expandable framework for the multi-modal sensing based on the partially paired sensing datasets, currently aligning six sensing modalities.
@@
  • We demonstrate Babel’s superior performance in modality enhancement and fusion. Additionally, we highlight Babel’s potential in the field of cross-modality retrieval, and its ability to bridge LLMs for enhanced comprehension of the physical world.
2. Background and Motivation
2.1. Multi-Modal Sensing
Sensing, through various modalities like vision, micro-electro-mechanical-system (MEMS) sensors, and wireless RF sensors, is now a ubiquitous method for comprehending the physical world, capturing a wide range of information from the environment or specific objects. Multiple sensing modalities provide complementary capabilities. For instance, LiDAR creates long-range 3D environmental maps for vehicles, ultrasonic sensors offer close-range detection, and cameras interpret road signs for driver-assistance systems. This combination leads to the concept of multi-modal sensing.
Efficient multi-modal sensing relies on effectively fusing insights from each modality. Early strategies used handcrafted heuristics to link modalities, like A3’s (Zhou et al., 2014) use of a gyroscope for smartphone attitude estimation, supplemented by calibration from magnetometer and accelerometer. However, these strategies aren’t scalable, especially with complex sensing signals and environments, making the construction of heuristics for each task increasingly unfeasible. Recently, DL has been used for multi-modal sensing, automatically revealing correlations among diverse sensing modalities through supervised or self-supervised learning, and show superior performance (Ihianle et al., 2020; Li et al., 2022; Radu et al., 2018; Chung et al., 2019; Salehi et al., 2022, 2022).
2.2. The Power of Modality Alignment
DL-based multi-modal sensing methods outperform heuristic ones due to modality alignment. The alignment restructures one modality’s feature according to the representation space of another, eventually projects each modality’s features into a unified representation space through the training. This space facilitates the ease of manipulation. Consequently, modality alignment offers a comprehensive understanding of sensing data by utilizing each modality’s unique strengths.
Beyond augmenting multi-modal sensing fusion, the modality alignment could further empower new sensing applications. As illustrated in Fig. 1, aligned modalities are mutually retrievable. Specifically, utilizing the joint feature space as a proxy, one could employ Wi-Fi channel state information (CSI) to derive the corresponding embedding in the visual modality, subsequently generating an image from the visual embeddings. Intuitively, this could be considered as an alternative realization of Wi-Fi imaging. The joint feature space also facilitates the use of multiple aligned modalities for retrieval, thereby enabling multi-modal imaging without necessitating additional intervention.
The rise of Large Language Models (LLMs), presents new opportunities for sensing to interact with and understand the physical world (Xu et al., 2024). Despite the development of multi-modal LLMs, the wide range of sensing modalities requires a unified sensing ontology for seamless integration with LLMs (Li et al., 2024). This alignment creates a unified representation that bridges the gap between various sensing modalities and LLMs.
To demonstrate the power of modality alignment, Babel aligns six prevalent sensing modalities, including vision, depth, IMU, Wi-Fi, mmWave, and LiDAR.
2.3. Challenges and Opportunities
Figure 2. Five public datasets, XRF55 (Wang et al., 2024), OPERANet (Bocus et al., 2021), MM-Fi (Yang et al., 2023c), UTD-MHAD (Chen et al., 2015) and Kinectics-400 (Kay et al., 2017) with paired data cover six sensing modalities.
@@ 2.3. Challenges and Opportunities
Modality alignment is a growing AI research field, involving various methods (van den Oord et al., 2018; Tian et al., 2019). Of these, contrastive learning (CL) is notable. CL, a self-supervised learning method, differentiates similar and dissimilar samples by comparing positive (similar) and negative (dissimilar) pairs. The goal is to generate representations where similar samples are close, and dissimilar ones are far apart in the feature space. Contrastive Language-Image Pretraining (CLIP) (Radford et al., 2021) exemplifies the effective use of CL in aligning text and image modalities. CLIP, trained on a large Internet image-caption pairs, learns to associate semantically related texts and images.
@@
Nonetheless, it is still challenging to apply CL to align multiple sensing modalities, due to the fundamental data scarcity issue. For instance, CLIP’s training necessitates approximately $400,000,000$ image-text pairs, a scale of data that public datasets with paired sensing samples fail to match. In fact, public multi-modal sensing datasets (Chen et al., 2015; An et al., 2022b; Chen et al., 2023), as an example, contain a mere $600-42,000$ sample pairs, a stark contrast to the required volume.
Additionally, there exists numerous sensing modalities. The alignment of N sensing modalities generally necessitates a substantial amount of N-tuple data. Regrettably, the most public dataset currently available usually contains triplets of paired sensing modalities. There are no public datasets that cater to the alignment of a greater number of sensing modalities, such as six or more.
Furthermore, the data scarcity issue cannot be easily resolved through large-scale data collection. Unlike visual or linguistic data, which are already available on the Internet, sensing data is generated at and confined to end clients. The collection of such data often necessitates specialized hardware or software. The transmission of this data is resource-intensive and poses significant privacy concerns, rendering ubiquitous sensing data collection impractical. The simultaneous collection of multi-modal sensing data even compounds this problem.
Thereby, the existing research (Ouyang et al., 2022; Yuan et al., 2024; Xu et al., 2023; Han et al., 2023; Moon et al., 2023; Girdhar et al., 2023) may struggle to align multiple sensing modalities in a scalable manner. For instance, ImageBind (Girdhar et al., 2023), and OneLLM (Han et al., 2023) are primarily constrained to supporting IMU due to the scarcity of paired data on other sensing modalities. Cosmo (Ouyang et al., 2022) can only align a limited array of sensing modalities as the demand for N-tupled paired data escalates when aligning N modalities. In Babel, we endeavor to align six sensing modalities with the existing constrained datasets, based on the following opportunities we identified.
Over the past few decades, researchers have developed a range of feature extractors or encoders for single modality sensing, based on either signal processing or DL techniques. These have been proven effective in extracting representative features for various downstream tasks. This presents an opportunity to leverage existing single modality encoders in constructing the modality alignment network. In this way, we might significantly reduce the trainable parameters, thereby decreasing the data requirement in accordance.
Furthermore, we observe that while no dataset provides N-tuple samples for N modalities, numerous datasets do contain paired data for two modalities. Despite these datasets potentially being collected for different tasks, they may share common modalities. Fig. 2 illustrates this occurrence, with five datasets encompassing six modalities. This presents an opportunity to assemble a compact dataset with restricted modalities and exploit the shared modalities to incrementally expand the modality alignment network.
To cope with the data scarcity challenge, and to capitalize on the aforementioned opportunities, we introduce Babel.

### v2

# Babel: A Scalable Pre-trained Model for Multi-Modal Sensing via Expandable Modality Alignment

@@ Abstract.
This paper presents Babel, the expandable modality alignment model, specially designed for multi-modal sensing. While there has been considerable work on multi-modality alignment, they all struggle to effectively incorporate multiple sensing modalities due to the data scarcity constraints. How to utilize multi-modal data with partial pairings in sensing remains an unresolved challenge.
Babel tackles this challenge by introducing the concept of expandable modality alignment. The key idea involves transforming the N-modality alignment into a series of binary-modality alignments. Novel techniques are also proposed to further mitigate data scarcity issue and balance the contribution of the newly incorporated modality with the previously established modality alignment during the expandable alignment process. We provide the comprehensive implementation. In the pre-training phase, Babel currently aligns 6 sensing modalities, namely Wi-Fi, mmWave, IMU, LiDAR, video, and depth. For the deployment phase, as a foundation model, any single or combination of aligned modalities could be selected from Babel and applied to downstream tasks.
Evaluation demonstrates Babel’s outstanding performance on eight human activity recognition datasets, compared to a broad range of baselines e.g., the SOTA single-modal sensing networks, multi-modal sensing framework, and multi-modal large language models. Babel not only improves the performance of individual modality sensing (12% averaged accuracy improvement), but also effectively fuses multiple available modalities (up to 22% accuracy increase). Case studies also highlight emerging application scenarios empowered by Babel, including cross-modality retrieval (i.e., sensing imaging), and bridging LLM for sensing comprehension. Our code is available at aka.ms/project-babel.
Multi-modal Sensing, Modality Alignment, Pre-trained Model, Human Activity Recognition
Research is done during internship at Microsoft Research
journalyear: 2025copyright: acmlicensedconference: The 23rd ACM Conference on Embedded Networked Sensor Systems; May 6-9, 2025; Irvine, USAccs: Computing methodologies Neural networksccs: Computing methodologies Activity recognition and understandingccs: Information systems Multimedia and multimodal retrievalccs: Hardware Sensor devices and platforms
@@ 1. Introduction
Sensing technology, with its distinctive capacity to perceive the physical world, has found widespread application across a multitude of domains, encompassing healthcare, mixed reality, smart driving, and beyond. Over the past several decades, a plethora of sensing modalities have been investigated, each offering a unique and complementary perspective of the physical world. This has led to the emergence of multi-modal sensing, an approach that harnesses the simultaneous use of multiple sensing modalities.
@@
Early methods for organizing multiple sensing modalities relied on handcrafted heuristics or features (Zhou et al., 2014), which is proved challenging to scale across various modalities and tasks, due to the complexity of sensing signals and environments. Recent advancements in multi-modal learning have introduced promising solutions (Zhao et al., 2018; Deldari et al., 2022; Jain et al., 2022). These methods automatically uncover correlations among diverse sensing modalities through supervised or self-supervised learning (Ihianle et al., 2020; Li et al., 2022; Radu et al., 2018; Chung et al., 2019; Salehi et al., 2022). Among them, modality alignment projects the representations of each sensing modality into a unified and shared space by leveraging paired modality data, demonstrating superior performance (Ouyang et al., 2022).
Although modality alignment can effectively organize sensing modalities, existing work (Ouyang et al., 2022; Radu et al., 2018; Chung et al., 2019; Salehi et al., 2022) is often tailored for specific modalities, necessitating resampling and retraining for different downstream tasks and modality combinations, hindering seamless deployment of sensing applications. Therefore, this paper poses the question: Can we build a pre-trained multi-modal sensing alignment network as a foundation model? This model would align common sensing modalities and allow for the integration of new modalities. In the deployment phase, any single or combination of aligned modalities from the model could be selected and applied to downstream tasks directly without retraining.
Figure 1. Five public sensing datasets, XRF55 (Wang et al., 2024), OPERANet (Bocus et al., 2021), MM-Fi (Yang et al., 2023c), UTD-MHAD (Chen et al., 2015) and Kinectics-400 (Kay et al., 2017) with binary-paired data cover six modalities.
@@
While modality alignment in AI is a growing research area, its application in sensing presents significant hurdles. The fundamental challenge in supporting multi-modality in sensing is the data scarcity, specifically, (i) the scarcity of paired data, which is essential for aligning two modalities. For instance, the widely-used CLIP (Radford et al., 2021) required 400 million image-text pairs for pre-training. In sensing, there lacks paired data from all modalities since some modality data require specialized hardware and expertise to collect. and (ii) the scarcity of multi-paired modalities. Existing datasets only contain data from a subset of modalities (Chen et al., 2015; Yang et al., 2023c; Bocus et al., 2021; Kay et al., 2017). For these reasons, existing research (Han et al., 2023; Yuan et al., 2024; Ouyang et al., 2022; Xu et al., 2023; Zhang et al., 2023a; Girdhar et al., 2023) struggle to fully incorporate multiple sensing modalities. For instance, due to the limited language-paired data, OneLLM (Han et al., 2023) supports a limited number of sensing modalities, i.e., IMU, with subpar performance (see Table 4). Cosmos (Ouyang et al., 2022) pioneered the alignment of multiple modalities, but due to the scarcity of multi-paired modalities, it aligns a limited number of modalities e.g., IMU and depth.
@@
In this paper, we present Babel to address this challenge, establishing the first scalable pre-trained network aligning multiple sensing modalities. The design of Babel is underpinned two observations: (i) Despite the scarcity of paired data, there exist well-developed encoders or feature extractors for single modality sensing. By leveraging these encoders, the amount of paired data required for modality alignment could be significantly reduced. (ii) Even though few datasets provide more than three paired modalities, numerous paired datasets exist that share common sensing modalities. These shared modalities can serve as a bridge for multi-modality alignment (see Fig. 1).
Drawing from these observations, the key idea of Babel is the expandable multi-modal alignment, which particularly transforms an N-modality alignment problem into a sequence of binary modality alignments. The expandability facilitates the effective utilization of partially paired data in the sensing community. As illustrated in Fig. 1, we could achieve alignment of six modalities through five binary-modality alignments, using the corresponding datasets.
@@
To realize this expandability, we introduce three techniques: the pre-trained modality tower (§4), the expandable network architecture (§5), and the adaptive training strategy (§6). Each modality utilizes a modality tower to extract features from raw data. We build these towers using existing singular-modal sensing feature extractors (e.g., LIMU-BERT (Xu et al., 2022) for IMU), and extent with our alignment modules for aligning with other modality towers; The expandable network architecture enables sequential training phases with only binary-paired samples. Within it we propose the prototype network, shared by all modalities, maintains the knowledge of aligned modalities when adding new ones. Lastly, our adaptive training strategy balances the contribution of newly added modalities to the unified representation, optimally assimilating new knowledge during model growth without disrupting established alignments.
@@
We offer a comprehensive implementation of Babel, including the network architecture, data preparation and processing, as well as the training details. In Babel, we currently align six common sensing modalities: two for wireless sensing, namely Wi-Fi and mmWave, two for mobile sensing, specifically IMU and LiDAR, and two for general vision, namely RGB and depth. As an expandable framework, Babel is allowed for aligning more modalities in the future without retraining aligned modalities. In our work, five datasets are utilized to construct Babel, including UTD-MHAD (Chen et al., 2015), Kinetics-400 (Kay et al., 2017), OPERANet (Bocus et al., 2021), XRF55 (Wang et al., 2024), and MM-Fi (Yang et al., 2023c).
The current pre-trained Babel is evaluated on a typical sensing application, Human Activity Recognition (HAR), across eight datasets, which include both in-domain and out-of-domain datasets (Chen et al., 2015; Yang et al., 2023c; Bocus et al., 2021; Wang et al., 2024; Reyes-Ortiz et al., 2016; Zhang et al., 2022; An et al., 2022a; Li et al., 2010). To demonstrate Babel’s capability, we compared it to an array of baselines, including the state-of-the-art (SOTA) singular-modal sensing networks (Xu et al., 2021; Yang et al., 2023a; Li et al., 2010; An et al., 2022b), multi-modal sensing framework (Ouyang et al., 2022), and the emerging Multi-modal Large Language Models (MLLMs) (Han et al., 2023; Yuan et al., 2024; Girdhar et al., 2023; Zhang et al., 2023a).
We use the one-shot learning [^1] For each HAR class, we just use only one sample to fine-tune the downstream task classification header. to evaluate Babel as a foundation model. The evaluation shows that, (i) owing to the pre-trained alignment across more sensing modalities and extensive datasets, Babel improve the accuracy of single modality sensing by up to 20%, and 12% on average on all six aligned modalities across various datasets. (ii) Thanks to the aligned unified embedding space, Babel increases multi-modal sensing fusion accuracy by up to 22% compared to current multi-modal frameworks (Ouyang et al., 2022). (iii) When comparing with emerging MLLMs, which are limited in the range of sensing modalities they support, Babel surpasses them by the accuracy of 25.2% across HAR datasets.
Besides HAR, we also present two application studies to highlight Babel’s potential as a foundation model. The first is sensing imaging to illustrate cross-modality retrieval. With Babel, the original image-to-image diffusion model can be supplemented with non-visual data as input to generate images (Ramesh et al., 2022). The other case aims to bridge the gap between LLM and sensing. By injecting the IMU sensing signal through Babel into the Video-LLaMA (Zhang et al., 2023b), the LLM can understand the sensing signals without any retraining.
@@
  • Babel, to the best of our knowledge, is the first expandable foundation model for the multi-modal sensing, currently aligning six sensing modalities.
@@
  • We demonstrate Babel’s superior performance compared to a range of baselines. Additionally, we highlight Babel’s potential in the field of cross-modality retrieval, and its ability to bridge LLMs for enhanced comprehension of the physical world.
2. Related Work
In this section, we introduce the work pertaining to the design of Babel, and highlight our distinctive contribution divergent from the existing research. Specifically, we discuss key advancements in modality alignment, multi-modal sensing, and multi-modal LLMs.
@@ 2. Related Work
Modality alignment, as an emerging research topic, involving various methods (van den Oord et al., 2018; Tian et al., 2019). Of these, contrastive learning (CL) is notable. CL, a self-supervised learning method, differentiates similar and dissimilar samples by comparing positive (similar) and negative (dissimilar) pairs. The goal is to generate representations where similar samples are close, and dissimilar ones are far apart in the feature space. Contrastive Language-Image Pretraining (CLIP) (Radford et al., 2021) exemplifies the effective use of CL in aligning text and image modalities. CLIP, trained on a large Internet image-caption pairs, learns to associate semantically related texts and images. FOCAL (Liu et al., 2024) introduces an innovative CL framework with a temporal structural constraint designed for sensing data, addressing challenges in capturing both shared and modality-exclusive features in multimodal time-series data.
@@
Nonetheless, it is still challenging to apply CL to align multiple sensing modalities, due to the data scarcity issue. For instance, CLIP’s training necessitates approximately $400,000,000$ image-text pairs, a scale of data that public datasets with paired sensing samples fail to match. In fact, public multi-modal sensing datasets (Chen et al., 2015; An et al., 2022b; Chen et al., 2023), as an example, contain a mere $600-42,000$ sample pairs, a stark contrast to the required volume. Additionally, there exists numerous sensing modalities. The alignment of N sensing modalities generally necessitates a substantial amount of N-tuple data. Regrettably, There are no public datasets that cater to the alignment of a greater number of sensing modalities, such as six or more. Babel addresses this fundamental challenge via the proposed expandable modality alignment technique.
Multi-modal sensing offers unique abilities to perceive the physical world, incorporates a plethora of methods. For instance, Cosmo (Ouyang et al., 2022) pioneered the application of contrastive fusion learning in multi-modal sensing, incorporating RGB, depth and IMU modalities. MESEN (Xu et al., 2023) employs multi-modal contrastive learning to improve the performance of singular-modal sensing. FM-Fi (Weng et al., 2024) leverages CLIP through cross-modal contrastive knowledge distillation to improve Radio-Frequency-based human activity recognition with limited labeled data. Nevertheless, these studies are typically crafted for chosen modalities, necessitating retraining for additional ones. In stark contrast to Cosmo, MESEN and FM-Fi, Babel operates as a pre-trained foundational model for multi-modal sensing, facilitating the utilization of any single or multiple aligned modalities for downstream sensing tasks without retraining. Moreover, due to the pre-trained alignment across a broad spectrum of modalities, Babel attains exceptional performance in both single-modal sensing and multi-modal fusion (§8).
The concept of multi-modal sensing is extensively employed across a wide array of applications. For instance, (Liu et al., 2021b) integrates RFID and RGB for recognizing human-object interactions. (Wang et al., 2022) leverages LiDARs, cameras, and IMU and GNSS devices worn by animals to recognize animal behavior. To locate target individuals, (Liu et al., 2021a) utilizes Wi-Fi Fine Timing Measurements and IMU data to associate individuals in a video with a matched query ID. GaitVibe+ (Dong et al., 2022) enhances structural vibration-based footstep localization using temporary cameras and vibration sensors for in-home gait analysis. (Girod and Estrin, 2001) presents an acoustic and camera sensing system that ameliorates range estimation for applications in robotics and others. These applications could benefit through Babel.
Multi-modal LLMs are rapidly evolving to accommodate an increasing number of modalities. Supporting an expanded range of modalities typically necessitates a multi-modal encoder, which projects multi-modal signals into the language embedding space. To construct such an encoder, Meta-Transformer (Zhang et al., 2023a) demonstrates the potential of employing a shared transformer encoder across 12 modalities. ImageBind (Girdhar et al., 2023) aligns six modalities utilizing solely image-paired data. Similarly, LanguageBind (Zhu et al., 2024; Zhou et al., 2023) employs the language as the central binding modality to align four modalities. OneLLM (Han et al., 2023) aligns eight modalities to language using a singular, unified encoder. CoDi (Tang et al., 2023) facilitates alignment across language, image, video, and audio modalities.
These studies, however, offer exceedingly limited support for sensing modalities. Indeed, the sole supported sensing modality is the IMU, yet its performance is notably subpar (see the detailed evaluation results in §8). This is primarily due to the high requirement for data with specific modality pairings, such as image-sensing pairs. For instance, ImageBind (Girdhar et al., 2023) is only trained on Ego4D (Grauman et al., 2022) for IMU. Their cross-domain capabilities are restricted and they cannot be trained on other sensing modalities without first addressing the issue of data scarcity. In response, we propose crucial techniques for aligning sensing modalities through an expandable architecture, reducing dependence on exhaustive modality pairings. Details are presented in the following.